OpenStudy (anonymous):

The sum of two consecutive even integers is 234. What is the larger integer?

hartnn (hartnn):

let the larger integer be x so smaller integer = x-2 sum=234 so x+x-2=234 so 2x=234+2 can u solve further?

OpenStudy (anonymous):

118?

hartnn (hartnn):

right :)

OpenStudy (anonymous):

ty :D

hartnn (hartnn):

welcome :)